Can you please explain how to multiply square roots? For example, √5(√5-5√3).

Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

√5(√5-5√3).
----
= 5 - 5*sqrt(15)
---------------------
Cheers,
Stan H.

Answer by lwsshak3(11628)

Can you please explain how to multiply square roots? For example, √5(√5-5√3)
**
radicans(numbers under the sqrt sign) can be multiplied and divided if the product or quotient remains under the sqrt sign. Perfect squares can be pulled out from under the sqrt sign.
√5(√5-5√3)=√25-5√15=5-5√15=5(1-√15)